1.    Given the following cost function: TC = 1500 + 15Q – 6Q 2 + Q3               

a.    Determine the total fixed cost for producing 1000 units of output and 500 units of output. 

b.    What is AFC at: 

                                              i.     1000 units of output b) 

                                            ii.     500 units of output iii. 

c.    Determine TVC, AVC, MC and AC at 50 units of output.



1
Expert's answer
2021-01-12T13:04:59-0500

a. Total fixed cost

at Q=1000

FC = 1500

at Q=500

FC = 1500

b. AFC "= \frac{FC}{Q}"

i. at Q = 1000

"AFC = \frac{1500}{1000} = 1.5"

ii. at Q = 500

"AFC = \frac{1500}{500}=3"

c. TVC = TC – FC = 15Q – 6Q2 + Q3

"AVC = \frac{TVC}{Q} = 15 - 6Q + Q^2 \\

AC = \frac{TC}{Q} = \frac{1500}{Q}+15-6Q+Q^2 \\

MC = \frac{dTC}{dQ}=15-12Q+3Q^2"

at Q = 50

"TVC = 15(50) -6(50)^2 + (50)^3 \\

= 750 -15000 + 125000 \\

= 110750 \\

AVC = \frac{TVC}{Q} \\

\frac{110750}{50}=2215 \\

AC = \frac{1500}{50}+15 - 6(50)+50^2 \\

= 2245 \\

MC = 15 -12(50) + 3(50)^2 \\

= 6915 \\

TVC = 110750 \\

AVC = 2215 \\

MC = 6915 \\

AC = 2245"